### v1

# Image captioning in different languages

Abstract
This short position paper provides a manually curated list of non-English image captioning datasets (as of May 2024). Through this list, we can observe the dearth of datasets in different languages: only 23 different languages are represented. With the addition of the Crossmodal-3600 dataset (Thapliyal et al., 2022, 36 languages) this number increases somewhat, but still this number is tiny compared to the thousands of spoken languages that exist. This paper closes with some open questions for the field of Vision & Language.
Image captioning in different languages
Emiel van Miltenburg Tilburg University C.W.J.vanMiltenburg@tilburguniversity.edu
1 Introduction
Image captioning datasets have been developed for several different languages, but only a handful (mostly English datasets) have received attention in the NLP community. This research note aims to provide an overview of the different image captioning datasets that have been described in the literature. This list is the result of monitoring Google Scholar alerts for the last eight years using the keywords "MS COCO" and "Flickr30K" (the prototypical English captioning datasets), and manually curating a list of papers that describe image captioning work in languages other than English.
1.1 Earlier work
An earlier version of this list appeared in the PhD thesis of van Miltenburg (2019). The list shows the language of the dataset, the source of the images (either an existing dataset such as Flickr8/30K and MS COCO or the name of the newly introduced dataset, whether the captions were translated from English or independently collected, and a reference to the paper where the dataset was introduced.
Next to individual references, I also relied on the survey by Jayaswal et al. (2023) to get an overview of image captioning work for Indian languages (i.e. Assamese, Bengali/Bangla, Hindi, Punjabi, Tamil, Telugu, and Urdu). Although the survey itself does not contain a list of existing datasets, the references provided a useful starting point to add all Indian captioning datasets to the list.
1.2 Additions
Given the creation history of this overview, I probably have overlooked some datasets. Let me know if you have developed a captioning dataset in a language other than English.
Language Source T I Citation
Arabic ArabicFashionData $\checkmark$ Al-Malki and Al-Aama 2023
Arabic MS COCO $\checkmark$ Mohamed et al. 2023
Arabic MS COCO* $\checkmark$ Mohamed et al. 2023
Arabic MS COCO* $\checkmark$ Al-muzaini et al. 2018
Arabic Flickr8K* $\checkmark$ Al-muzaini et al. 2018
Arabic MS COCO* $\checkmark$ Al-muzaini et al. 2018
Assamese Assamese-news $\checkmark$ Das and Singh 2021, 2022
Assamese Flickr30K $\checkmark$ Nath et al. 2022
Assamese MS COCO $\checkmark$ Nath et al. 2022
Bengali BanglaLekha $\checkmark$ Rahman et al. 2019
Bengali Bornon $\checkmark$ Muhammad Shah et al. 2021
Bengali Flickr8K $\checkmark$ Muhammad Shah et al. 2021; Humaira et al. 2021
Bengali BNLIT $\checkmark$ Jishan et al. 2020
Bengali Flickr30K $\checkmark$ Nawer et al. 2023
Brazilian Portuguese Flickr30K $\checkmark$ $\checkmark$ Viridiano et al. 2024
Chinese Flickr8K $\checkmark$ $\checkmark$ Li et al. 2016
Chinese Flickr30K $\checkmark$ $\checkmark$ Lan et al. 2017
Chinese MS COCO* $\checkmark$ $\checkmark$ Li et al. 2018
Chinese AI Challenger/ICC $\checkmark$ Wu et al. 2017
Chinese Thangka $\checkmark$ Hu et al. 2023
Czech Flickr30K $\checkmark$ Barrault et al. 2018
Dutch Flickr30K* $\checkmark$ van Miltenburg et al. 2017
Dutch MS COCO* $\checkmark$ van Miltenburg et al. 2018
French MS COCO* $\checkmark$ Rajendran et al. 2016
French Flickr30K $\checkmark$ Elliott et al. 2017
German MS COCO* $\checkmark$ Rajendran et al. 2016
German MS COCO* $\checkmark$ Hitschler et al. 2016
German Flickr30K $\checkmark$ $\checkmark$ Elliott et al. 2016
German IAPR-TC12 $\checkmark$ Grubinger et al. 2006
Hindi Flickr8K $\checkmark$ Rathi 2020
Hindi MS COCO $\checkmark$ Dhir et al. 2019; Mishra et al. 2021a, b
Hindi Visual Genome $\checkmark$ Parida et al. 2019
Indonesian Flickr8K $\checkmark$ Wijadi et al. 2024
Indonesian MS COCO $\checkmark$ Al-Faruq and Fudholi 2023
Japanese UIUC Pascal $\checkmark$ Funaki and Nakayama 2015
Japanese MS COCO* $\checkmark$ Miyazaki and Shimizu 2016
Japanese MS COCO $\checkmark$ Yoshikawa et al. 2017
Persian ParsVQA-Caps $\checkmark$ Mobasher et al. 2022
Portuguese #PraCegoVer $\checkmark$ Santos et al. 2021
Punjabi Punjabi Tribune $\checkmark$ Kaur et al. 2021
Spanish IAPR-TC12 $\checkmark$ Grubinger et al. 2006
Tamil MS COCO $\checkmark$ Vishnu Kumar and Lalithamani 2022
Tamil MS COCO $\checkmark$ B et al. 2023
Tamil Flickr8K $\checkmark$ B et al. 2023
Tamil Flickr30K $\checkmark$ B et al. 2023
Telugu MS COCO $\checkmark$ B et al. 2023
Telugu Flickr8K $\checkmark$ B et al. 2023
Telugu Flickr30K $\checkmark$ B et al. 2023
Telugu Flickr8K $\checkmark$ Dwarampudi et al. 2021
Thai HouseDefects $\checkmark$ Jaruschaimongkol et al. 2023
Thai Thai-Food $\checkmark$ Jaknamon and Marukatat 2022
Thai Thai-Travel $\checkmark$ Jaknamon and Marukatat 2022
Thai Flickr30K $\checkmark$ Jaknamon and Marukatat 2022
Turkish Flickr8K $\checkmark$ Unal et al. 2016
Turkish MS COCO $\checkmark$ Golech et al. 2022
Turkish MS COCO $\checkmark$ Yıldız et al. 2023
Ukrainian Flickr30K $\checkmark$ Saichyshyna et al. 2023
Urdu Flickr8K* $\checkmark$ Ilahi et al. 2020
Urdu Flickr8K* $\checkmark$ Afzal et al. 2023
Vietnamese MS COCO* $\checkmark$ Lam et al. 2020
Vietnamese KTVIC $\checkmark$ Pham et al. 2024
Vietnamese UIT-OpenViIC $\checkmark$ Bui et al. 2023
Table 1: Image description datasets available in languages other than English, with an indication of their source, and whether the descriptions were Translated or Independently collected. Asterisks indicate that the data is a subset of the original dataset. Flickr8K is the predecessor of Flickr30K, see Hodosh et al. 2013. For some recources multiple references are given since it is not clear which is the ‘main’ reference.
2 The list
Table 1 provides a list of non-English captioning datasets. [^1] The Crossmodal-3600 dataset (Thapliyal et al., 2022, 36 languages) is not listed in the table because it would result in too much repetition., [^2] Some authors opt not to translate the dataset but instead use an English-language captioning model and then translate the results (e.g. Tien et al. 2020). These works are also not in the table. Some observations:
  • Starting with the most obvious observation: although I made an effort to be inclusive and bookmark all non-English captioning datasets I came across, the table is still not very diverse. Only a handful of languages have researchers actively producing image captioning datasets.
  • Many datasets have been created through (automatic) translation of English captions, which is relatively cheap compared to independently collected captions. The downside, of course, is that these datasets still capture a Western perspective, even if the dataset has been written in a non-Western language.
  • Although Table 1 differentiates between translated and independently collected descriptions, in practice this distinction may be blurred by annotators providing manual corrections of automatic translations, sometimes replacing the translations altogether.
  • Some datasets are very specific to a particular use case, for example the Arabic Fashion dataset Al-Malki and Al-Aama (2023), the Chinese Thangka dataset Hu et al. (2023), and the Thai house defects dataset Jaruschaimongkol et al. (2023). These may not be useful for general image captioning tasks, but may still help to ground multi-modal language models.
  • Most of the papers in Table 1 have not been published at ACL workshops (or their Computer Vision equivalents), meaning that these are easily missed if you aren’t explicitly looking for them.
3 Research questions
Browsing the list above, many research questions come to mind. For example:
Assessment
How well do existing Vision & Language models perform on different datasets, such as Assamese-news, BNLIT, or Thai-Food/Travel? And, since any caption task inherently makes (implicit) assumptions about the intended audience, from which perspective should we assess these models?
Benefiting from multi-lingual data
How can we benefit from multi-lingual datasets? And how do we overcome the challenges brought by the imbalance in available training data?
Inclusion and Community Building
How inclusive are Vision & Language research and citation practices? Are we doing enough to accommodate and support speakers of underrepresented languages?
The last point clearly stands out from the previous two, in that it does not only require desk research but also carrying out community work, talking to different stakeholders, and bridging the gap between different cultures. This seems most important to make Vision & Language models work for everyone, yet this work is not strongly incentivized.
4 Conclusion
This note provides an overview of non-English captioning datasets. This is not only useful for those who would like to train a captioning model for the different languages that are listed in the table, but it hopefully also increases the visibility of the authors who have been working to promote image captioning technology for their own languages. It is also clear that much work remains to be done, not only developing datasets and models for different languages and cultures, but also in terms of building a community that embraces cultural diversity.

### v2

@@
 Abstract
-This short position paper provides a manually curated list of non-English image captioning datasets (as of May 2024).
-Through this list, we can observe the dearth of datasets in different languages: only 23 different languages are represented.
-With the addition of the Crossmodal-3600 dataset (Thapliyal et al., 2022, 36 languages) this number increases somewhat, but still this number is tiny compared to the thousands of spoken languages that exist.
-This paper closes with some open questions for the field of Vision & Language.
+[…]
 Image captioning in different languages
@@
 Arabic ArabicFashionData $\checkmark$ Al-Malki and Al-Aama 2023
+Arabic MS COCO* $\checkmark$ ElJundi et al. 2020
 Arabic MS COCO $\checkmark$ Mohamed et al. 2023
@@
 Arabic Flickr8K* $\checkmark$ Al-muzaini et al. 2018
-Arabic MS COCO* $\checkmark$ Al-muzaini et al. 2018
 Assamese Assamese-news $\checkmark$ Das and Singh 2021, 2022
@@
   • Most of the papers in Table 1 have not been published at ACL workshops (or their Computer Vision equivalents), meaning that these are easily missed if you aren’t explicitly looking for them.
+  • There are six languages in this list that are not in the Crossmodal-3600 dataset.
+These are: Assamese, Brazilian Portuguese, Portuguese, Punjabi, Tamil, and Urdu.
+This brings the total number of languages with available caption data to 36 + 6 = 42.
 3 Research questions
@@
 This seems most important to make Vision & Language models work for everyone, yet this work is not strongly incentivized.
+4 Discussion
+It is often said that there are about 7,000 languages in the world, which may lead us to think that we should also develop NLP technology for all of those languages.
+But Bird (2024) reminds us that beyond the 500 institutional languages, there are 6,500 oral languages that may not require technological solutions such as image captioning.
+[^3] See Bird and Yibarbuk (2024) for a more in-depth discussion of these numbers.
+Instead, the contact languages used between ‘outsiders’ and speakers of those oral languages (that overlap with the 500 institutional languages) should be the target language for NLP technology.
 5 Conclusion

### v3

@@
 Language Source T I Citation
+Arabic JEEM $\checkmark$ Kadaoui et al. 2025
 Arabic ArabicFashionData $\checkmark$ Al-Malki and Al-Aama 2023
@@
 Assamese Flickr30K $\checkmark$ Nath et al. 2022
+Assamese $\hookrightarrow$ revised $\checkmark$ Choudhury et al. 2023
 Assamese MS COCO $\checkmark$ Nath et al. 2022
+Assamese $\hookrightarrow$ revised $\checkmark$ Choudhury et al. 2023
 Bengali BanglaLekha $\checkmark$ Rahman et al. 2019
@@
 Bengali Flickr8K $\checkmark$ Muhammad Shah et al. 2021; Humaira et al. 2021
+Bengali Flickr8K $\checkmark$ Khan et al. 2022
 Bengali BNLIT $\checkmark$ Jishan et al. 2020
 Bengali Flickr30K $\checkmark$ Nawer et al. 2023
+Bengali Flickr30K $\checkmark$ Hossain and Hasan 2023
 Brazilian Portuguese Flickr30K $\checkmark$ $\checkmark$ Viridiano et al. 2024
@@
 Spanish IAPR-TC12 $\checkmark$ Grubinger et al. 2006
+Table 1: Image description datasets available in languages other than English (A-S), with an indication of their source, and whether the descriptions were Translated or Independently collected.
+Asterisks indicate that the data is a subset of the original dataset.
+Flickr8K is the predecessor of Flickr30K, see Hodosh et al.
+2013.
+For some recources multiple references are given since it is not clear which is the ‘main’ reference.
+In the case of Assamese, Choudhury et al.
+(2023) manually corrected translation errors from the datasets provided by Nath et al.
+(2022), resulting in a revised (but highly similar) dataset.
+Language Source T I Citation
 Tamil MS COCO $\checkmark$ Vishnu Kumar and Lalithamani 2022
@@
 Turkish MS COCO $\checkmark$ Yıldız et al. 2023
+Turkish Tiny TR-cap $\checkmark$ Memiş and Yıldız 2025
 Ukrainian Flickr30K $\checkmark$ Saichyshyna et al. 2023
@@
 Vietnamese UIT-OpenViIC $\checkmark$ Bui et al. 2023
-Table 1: Image description datasets available in languages other than English, with an indication of their source, and whether the descriptions were Translated or Independently collected.
+Table 2: Image description datasets available in languages other than English (T-Z), with an indication of their source, and whether the descriptions were Translated or Independently collected.
 Asterisks indicate that the data is a subset of the original dataset.
@@
 2 The list
-Table 1 provides a list of non-English captioning datasets.
+Tables 1 and 2 provide a list of non-English captioning datasets.
 [^1] The Crossmodal-3600 dataset (Thapliyal et al., 2022, 36 languages) is not listed in the table because it would result in too much repetition., [^2] Some authors opt not to translate the dataset but instead use an English-language captioning model and then translate the results (e.g.
@@
 The downside, of course, is that these datasets still capture a Western perspective, even if the dataset has been written in a non-Western language.
-  • Although Table 1 differentiates between translated and independently collected descriptions, in practice this distinction may be blurred by annotators providing manual corrections of automatic translations, sometimes replacing the translations altogether.
+  • Although Tables 1 and 2 differentiate between translated and independently collected descriptions, in practice this distinction may be blurred by annotators providing manual corrections of automatic translations, sometimes replacing the translations altogether.
   • Some datasets are very specific to a particular use case, for example the Arabic Fashion dataset Al-Malki and Al-Aama (2023), the Chinese Thangka dataset Hu et al.